OpenStudy (anonymous):

any body know how to integrate (1)/(x^3+3x^2) using partial fractions?

OpenStudy (anonymous):

First step, factor out the bottom. You'll get \(x^2(x+3)\). Then, set it up like this: \[ \frac{A}{x}+\frac{B}{x^2}+\frac{C}{x+3} \]Set that equal to your equation, and multiply everything by the denominator.\[ 1=A(x^2+3x)+B(x+3)+C(x^2)\\ 1=(A+C)x^2+(3A+B)x+3B \]Now you have the following system of equations:\[ A+C=0\\ 3A+B=0\\ 3B=1 \]This means that:\[ B=\frac{1}{3}, A=\frac{-1}{9}, C=\frac{1}{9} \]So, the equation can be rewritten as:\[ \frac{\frac{-1}{9}}{x}+\frac{\frac{1}{3}}{x^2}+\frac{\frac{1}{9}}{x+3} \]You can test that this is right by recombining the fractions to see if it adds back up correctly. Now, integrate each term individually:\[ \frac{-1}{9}\int\frac{1}{x}dx+\frac{1}{3}\int\frac{1}{x^2}dx+\frac{1}{9}\int\frac{1}{x+3}dx \]

OpenStudy (anonymous):

shouldn't it be A(x^2)(x+3)+B(x)(x+3)+C(x)(x^3)?

OpenStudy (anonymous):

Nope, that's a common misconception. You don't multiply by each of the other denominators. You multiply the whole thing by the original denominator.

OpenStudy (anonymous):

sorry. still not understanding that part

OpenStudy (anonymous):

You don't need to multiply A by \(x^2(x+3)\), because A is already over an x, so you just need \(x(x+3)\) to reach the common denominator.

OpenStudy (anonymous):

\[ \frac{1}{x^2(x+3)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x+3}\\ \frac{x^2(x+3)}{x^2(x+3)}=\frac{Ax^2(x+3)}{x}+\frac{Bx^2(x+3)}{x^2}+\frac{Cx^2(x+3)}{x+3}\\ 1=Ax(x+3)+B(x+3)+C(x^2) \]

OpenStudy (anonymous):

I would use Heavy side cover up for B and C and for A just one substitution.

OpenStudy (anonymous):

thanks guys. i got it. appreciate it!

OpenStudy (anonymous):

since we know what B and C is, how would find A? I set x=-3 to find C. To find B, i set x=0.

OpenStudy (anonymous):

My method is just to treat it as a system of equations, as above, and solve the system.

OpenStudy (anonymous):

this is a different question. how would you separate 4x+16/(x(x-2)^2) into the A, B, and C form?

OpenStudy (anonymous):

Please close this question and open a new one to ask a new question :)